# Text-to-Decision Agent: Offline Meta-Reinforcement Learning from Natural Language Supervision

### v4

@@ 1 Introduction
Reinforcement learning (RL) has emerged as an effective mechanism for training autonomous agents to perform complex tasks in interactive environments [1, 2], unleashing its potential across frontier problems including preference optimization [3], diffusion model training [4], and reasoning such as in OpenAI o1 [5] and DeepSeek-R1 [6]. Despite its achievements, one of the grand challenges is generalization: building a general-purpose agent capable of handling multiple tasks in response to diverse user commands [7]. Due to distribution shift and the lack of self-supervised pre-training techniques, RL agents typically struggle with poor generalization to unseen tasks [8]. Offline meta-RL tackles generalization via training on a distribution of offline tasks [9]. However, they usually rely on high-quality samples [10] or warmup explorations [11] to infer task beliefs for generalization at test time, while these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks [12]. This inspires us to explore a much broader source of supervision.
@@
Recently, large language models (LLMs), pre-trained on extensive text corpora, can encode a wealth of semantic knowledge with remarkable representation power and transferability [13]. The typical “text-to-text" interface enables task-agnostic architectures to achieve broad generalization with minimal or no domain-specific data [14]. The availability of large-scale text collections for aggregate supervision during model pre-training has revolutionized the language and multi-modal communities in recent years [15]. Naturally, the above advancements highlight a promising question: Could scalable pre-training methods that learn perception from the supervision embedded in natural language lead to a similar leap forward in developing generalist agents for offline meta-RL?
@@
Figure 1: t-SNE visualization of Ant-Dir where tasks with target directions in $[0,2\pi]$ are mapped to rainbow-colored points. Top: we encode multi-task data into dynamics-aware decision embeddings to capture task-specific environment dynamics. Bottom: we bridge the semantic gap between text and decision via contrastive pre-training. The aligned text embeddings follow a cyclic spectrum that exactly matches the periodicity of angular directions in a physical sense. This interesting finding shows that we effectively align text embeddings to comprehend environment dynamics and facilitate convincing language grounding in decision domains.
@@
RL typically learns from active interactions with the outer world tied to specific environment dynamics, with unique decision-level representation distinct from the unbounded perception-level representation of natural language [16]. Recent studies facilitate language grounding to decision domains from various aspects, such as LLMs as policies [17, 18], LLMs as rewards [19, 20], and language-conditioned policy learning [21, 22]. Despite these efforts, how to harness unbounded representations of natural language knowledge for building generalist decision agents remains the following challenges: i) LLMs, trained on text corpora, typically lack grounding in the physical world and fail to capture any environment dynamics; ii) Leveraging LLMs for decision-making is prone to knowledge misalignment due to the semantic gap between the text and decision modalities; iii) Scalable implementations are necessary to fully harness language knowledge to train generalist decision models.
@@
To tackle these challenges, we propose Text-to-Decision Agent (T2DA), a simple and scalable pre-training framework for offline meta-RL via aligning language knowledge with environment dynamics of decision tasks. First, we pre-train a generalized world model to encode multi-task data into dynamics-aware decision embeddings, effectively capturing task-specific environment dynamics. Second, inspired by CLIP [14], we predict which textual description goes with which decision embedding, efficiently bridging their semantic gap via contrastive language-decision pre-training. It distills the world model structure to the text modality and aligns text embeddings to comprehend the environment dynamics. Finally, we deploy T2DA on two mainstream conditional generation architectures, training generalist policies conditioned on aligned text embeddings. During evaluation, natural language is used to reference learned decision perceptions or describe new ones, and the agent can directly realize zero-shot text-to-decision generation according to textual instructions at hand. Extensive experiments show that T2DA transfers nontrivially to downstream tasks, achieves high-capacity zero-shot generalization, and outperforms various types of baselines.
@@ 2 Related Work
Offline Meta-RL learns adaptable policies from pre-collected datasets without requiring online environment interaction [9, 23]. It addresses two fundamental challenges: the distributional shift between behavior and learned policies in offline RL [24, 25], and the quick adaptation to unseen tasks with minimal data. Existing approaches can be broadly categorized into optimization-based methods like MAML [26] and MACAW [9], and context-based methods such as PEARL [27], VariBAD [28], CORRO [23], CSRO [11], and UNICORN [29]. Recent studies of in-context RL attempt to leverage the in-context learning capability of transformers to improve RL’s generalization via casting RL as an across-episodic sequential prediction problem with few-shot prompting at test time [30, 31], such as AD [32] and DPT [33]. In general, the above approaches rely on high-quality samples or domain knowledge to infer task beliefs during evaluation, limiting their generality with this restricted form of supervision. This inspires us to explore a much broader source of supervision from natural language.
@@
LLMs for RL. Existing studies on leveraging LLMs for RL domains can be categorized into three main threads: LLMs as rewards, LLMs as policies, and language-conditioned policy learning. The first kind uses LLMs or VLMs (vision-language models) to automate the generation and shaping of reward functions [34, 19, 35, 20] or state representations [36], providing denser information to facilitate RL training. The second kind utilizes LLMs or VLMs as the policy backbone [16, 17, 37, 18]. Vision-language-action (VLA) models provide a direct instantiation of using pre-trained VLMs for robotics, fine-tuning visually-conditioned VLMs to generate robot control actions, such as RT-2 [38], OpenVLA [39], and $\pi_{0.5}$ [40]. In contrast, we leverage LLMs to obtain task representations for offline meta-RL and train conventional RL policies based on decision transformer or diffuser architectures.
@@
Language-conditioned policy. We focus on the third kind that learns a language-conditioned policy, extracting the world knowledge encoded in LLMs to help train offline meta-RL agents. Previous studies acquire language embeddings using pre-trained LLMs like BERT and GPT [41], while these embeddings are learned independently from decision tasks and may fail to capture domain-related information. Early methods develop rule-based [42] or task-specific [43] intermediate representations to capture task-related semantics, which can require a laborious design and lack scalability. Under an imitation learning paradigm, BC-Z [21] trains a vision-based robotic manipulation policy conditioned on pre-trained language embeddings, and BAKU [44] produces a simple transformer architecture that fuses multi-modal vision-language information and temporal context for multi-task decision-making. To improve language grounding, [45] translates natural language to task language with a referential game, and [22] uses supervised fine-tuning to enable bidirectional translation between language-described skills and rollout data. Overall, the challenges include the inability to capture environment dynamics, knowledge misalignment due to semantic gaps, and limited model scalability. In the paper, we tackle these challenges to fully harness LLMs for building generalist decision agents.
@@ 3 Method
In this section, we present Text-to-Decision Agent (T2DA), a simple and scalable framework to train generalist policies via effectively grounding natural language supervision with environment dynamics of decision tasks. Figure 2 illustrates the overall pipeline, followed by detailed implementations in the subsequent subsections. Corresponding algorithm pseudocodes are given in Appendix A.
@@ 3.3 Contrastive Language-Decision Pre-training
To successfully instruct generalist training with human language, we must bridge the semantic gap between text and the “decision modality". Inspired by CLIP [14], we predict which textual description goes with which training task, bridging their semantic gap via contrastive learning. The core idea is to connect the language supervision with corresponding decision embeddings, distilling the world model structure to the text modality and aligning text embeddings to comprehend environment dynamics.
Formally, given a batch of $N$ (trajectory $\tau$, textual task description $l$) pairs, we predict which of the $N\times N$ possible (trajectory, text) pairings across a batch actually occurred. We use the pre-trained trajectory encoder in Sec. 3.2 to extract the dynamics-aware decision embedding from the trajectory as $z=\phi(\tau)$. The text encoder is a transformer-based model $\psi$ that could be initialized from one of a wide variety of popular language models, such as CLIP [14] or T5 [46]. Then, the text embedding $z_{T}$ is easily derived by tokenizing the textual instruction $l$ and feeding it to the encoder as $z_{T}=\psi(l)$.
We model the pre-training task as a text-decision multi-modal learning problem by fine-tuning the text encoder while keeping the trajectory encoder fixed. The objective is to maximize the cosine similarity of the decision and text embeddings of the $N$ real pairs in the batch while minimizing the cosine similarity of the embeddings of the $N^{2}-N$ incorrect pairings. The symmetric similarities between the two modalities measured by cosine distance are defined as
$\text{sim}(\tau,l)\!=\!\exp(\alpha)\cdot\frac{\phi(\tau)W_{D}\cdot\psi(l)W_{T}}{||\phi(\tau)W_{D}||\cdot||\psi(l)W_{T}||},~{}~{}\text{sim}(l,\tau)\!=\!\exp(\alpha)\cdot\frac{\psi(l)W_{T}\cdot\phi(\tau)W_{D}}{||\psi(l)W_{T}||\cdot||\phi(\tau)W_{D}||},$ (3)
@@
$p(\tau_{k})=\frac{e^{\text{sim}(\tau_{k},l_{k})}}{\sum_{i=1}^{N}e^{\text{sim}(\tau_{i},l_{i})}},~{}~{}~{}p(l_{k})=\frac{e^{\text{sim}(l_{k},\tau_{k})}}{\sum_{i=1}^{N}e^{\text{sim}(l_{i},\tau_{i})}}.$ (4)
@@
where CE is the cross-entropy between two distributions. To effectively adapt the text encoder to the semantics of decision modality at a lightweight computing cost, we employ low-rank adaptation (LoRA) [47] to fine-tune the text encoder with the above contrastive loss. Experiments in Appendix E present the analysis of its efficiency and superiority compared to full-parameter fine-tuning.
@@ 3.4 Generalist Policy Learning
A widely adopted manner to tackle generalization in RL domains is to condition the policy on some representation that can capture task-relevant information, such as context-based meta-learning [27, 28] and prompt-based methods [10, 12]. Under this unified framework, we formulate the generalist agent as a task-conditioned policy model. There exists a true variable that represents the task identity, and we use a latent representation $h$ to approximate that variable. Then, a base policy $\pi(\cdot,h)$ can be shared across tasks by conditioning on the task representation as $\pi_{k}(a|s)\!=\!\pi(a|s;h_{k}),\forall k$. Previous studies usually approximate the task representation by acquiring expert decision data [10] or exploring the unseen task first [11]. This restricted form of supervision limits their universality and practicability since expensive high-quality samples or explorations are needed to specify any new task.
@@
3.5 Scalale Implementations
We develop scalable implementations of T2DA using two mainstream conditional generation architectures that hold the promise to train RL models at scale: the decision diffuser [48] and the decision transformer [49], yielding the following T2DA variants.
Text-to-Decision Diffuser (T2DA-D). Inspired by the recent success of text-to-image diffusion models [50, 51], we implement the text-to-decision counterpart by modeling the policy as a return-conditioned diffusion model with action planning. For model simplicity and scalability, we choose to diffuse over the $H$-step state-action trajectory as
@@ 3.5 Scalale Implementations
where $c$ denotes the timestep in the forward diffusion process, and $c\!=\!0$ corresponds to the unperturbed data.  [^1] Decision diffuser diffuses over only states and trains another inverse dynamics model to generate actions. In contrast, we only keep a unified diffusion model to i) ensure simplicity with a single end-to-end model in a multi-task setting, and ii) unlock the potential of diffusion models at scale, rather than depending on additional MLPs to generate part of the information, e.g., actions. To enable a generalist policy with natural language supervision, we include the fine-tuned text embedding $\psi(l)$ as an additional condition to the diffusion model. To this end, the generalist policy training is formulated as the standard problem of conditional generative modeling as
@@
Text-to-Decision Transformer (T2DA-T). We also implement T2DA with the causal transformer architecture, drawing upon the simplicity, scalability, and associated advances in language modeling such as BERT and GPT. We prepend a task prompt to the input of the causal transformer to realize generalization across tasks, akin to Prompt-DT [10]. For each task, we take the prompt-augmented trajectory $\tau^{+}$ as the input, which contains both the fine-tuned text embeddings $\psi(l)$ and the most recent $H$-step history sampled from offline datasets as
@@ 4 Experiments
Environments. We evaluate T2DA on three benchmarks that are widely adopted to assess generalization capacities of RL algorithms: i) the 2D navigation Point-Robot; ii) the multi-task MuJoCo locomotion control, containing Cheetah-Vel and Ant-Dir; and iii) the Meta-World platform for robotic manipulation. For each domain, we randomly sample a distribution of tasks captioned with text descriptions, and split them into training and test sets. We employ SAC [52] to independently train a single-task policy on each training task for offline dataset collection. We develop three types of offline datasets: Mixed, Medium, and Expert, following the common practice in offline RL [53]. Appendix B presents more details of environments and datasets.
Baselines. We compare T2DA to five competitive baselines that cover three representative paradigms in tackling RL generalization: the context-based offline meta-RL approaches, including 1) CSRO [11] and 2) UNICORN [29]; the in-context RL method of 3) AD [32]; and the language-conditioned policy learning methods including 4) BC-Z [21] and 5) BAKU [44].
For evaluation, all experiments are conducted using five different random seeds. The mean of received return is plotted as the bold line, with 95% bootstrapped confidence intervals of the mean indicated by the shaded region. Appendix D gives detailed implementations of T2DA and Appendix E gives analysis on fine-tuning of the text encoder.

### v5

@@ 1 Introduction
Reinforcement learning (RL) has emerged as an effective mechanism for training autonomous agents to perform complex tasks in interactive environments (mnih2015human; fawzi2022discovering), unleashing its potential across frontier problems including preference optimization (rafailov2023direct), diffusion model training (black2024training), and reasoning (cui2025entropy; yan2025learning) such as in OpenAI o1 (jaech2024o1) and DeepSeek-R1 (guo2025deepseek). Despite its achievements, one of the grand challenges is generalization: building a general-purpose agent capable of handling multiple tasks in response to diverse user commands (black2024pi). Due to distribution shift and the lack of self-supervised pre-training techniques, RL agents typically struggle with poor generalization to unseen tasks (schmied2023learning). Offline meta-RL tackles generalization via training on a distribution of offline tasks (mitchell2021offline). However, they usually rely on high-quality samples (xu2022prompting) or warmup explorations (gao2023context) to infer task beliefs for generalization at test time, while these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks (wang2024meta). This inspires us to explore a much broader source of supervision.
@@
Recently, large language models (LLMs), pre-trained on extensive text corpora, can encode a wealth of semantic knowledge with remarkable representation power and transferability (achiam2023gpt). The typical “text-to-text" interface enables task-agnostic architectures to achieve broad generalization with minimal or no domain-specific data (radford2021learning). The availability of large-scale text collections for aggregate supervision during model pre-training has revolutionized the language and multi-modal communities in recent years (schlarmann2024robust). Naturally, the above advancements highlight a promising question: Could scalable pre-training methods that learn perception from the supervision embedded in natural language lead to a similar leap forward in developing generalist agents for offline meta-RL?
@@
Figure 1: t-SNE visualization of Ant-Dir where tasks with target directions in $0,2\pi]$ are mapped to rainbow-colored points. Top: we encode multi-task data into dynamics-aware decision embeddings to capture task-specific environment dynamics. Bottom: we bridge the semantic gap between text and decision via contrastive pre-training. The aligned text embeddings follow a cyclic spectrum that exactly matches the periodicity of angular directions in a physical sense. This interesting finding shows that we effectively align text embeddings to comprehend environment dynamics and facilitate convincing language grounding in decision domains.
@@
RL typically learns from active interactions with the outer world tied to specific environment dynamics, with unique decision-level representation distinct from the unbounded perception-level representation of natural language (ahn2022can). Recent studies facilitate language grounding to decision domains from various aspects, such as LLMs as policies (szot2024large; hu2025divide; zhan2025exgrpo), LLMs as rewards (rocamonde2024vision; xie2024text2reward; fu2024furl), and language-conditioned policy learning (jang2022bc; pang2024knowledgeable; haldar2024baku). Despite these efforts, how to harness unbounded representations of natural language knowledge for building generalist decision agents remains the following challenges: i) LLMs, trained on text corpora, typically lack grounding in the physical world and fail to capture any environment dynamics; ii) Leveraging LLMs for decision-making is prone to knowledge misalignment due to the semantic gap between the text and decision modalities; iii) Scalable implementations are necessary to fully harness language knowledge to train generalist decision models.
@@
To tackle these challenges, we propose Text-to-Decision Agent (T2DA), a simple and scalable pre-training framework for offline meta-RL via aligning language knowledge with environment dynamics of decision tasks. First, we pre-train a generalized world model to encode multi-task data into dynamics-aware decision embeddings, effectively capturing task-specific environment dynamics. Second, inspired by CLIP (radford2021learning), we predict which textual description goes with which decision embedding, efficiently bridging their semantic gap via contrastive language-decision pre-training. It distills the world model structure to the text modality and aligns text embeddings to comprehend the environment dynamics. Finally, we deploy T2DA on two mainstream conditional generation architectures, training generalist policies conditioned on aligned text embeddings. During evaluation, natural language is used to reference learned decision perceptions or describe new ones, and the agent can directly realize zero-shot text-to-decision generation according to textual instructions at hand. Extensive experiments show that T2DA transfers nontrivially to downstream tasks, achieves high-capacity zero-shot generalization, and outperforms various types of baselines.
@@ 2 Related Work
Offline Meta-RL learns adaptable policies from pre-collected datasets without requiring online environment interaction (mitchell2021offline; yuan2022robust). It addresses two fundamental challenges: the distributional shift between behavior and learned policies in offline RL (levine2020offline; kumar2020conservative), and the quick adaptation to unseen tasks with minimal data. Existing approaches can be broadly categorized into the memory-based (e.g., RL2 (duan2016rl) and LLIRL (wang2022lifelong; wang2023dirichlet)), the optimization-based (MAML (finn2017model) and MACAW (mitchell2021offline)), and the context-based methods (e.g., PEARL (rakelly2019efficient), VariBAD (zintgraf2021varibad), etc. (yuan2022robust; gao2023context; li2024towards)). Recent studies of in-context RL attempt to leverage the in-context learning capability of transformers to improve RL’s generalization via casting RL as an across-episodic sequential prediction problem with few-shot prompting at test time (huang2024incontext; grigsby2024amago; wu2025mixture; liu2025scalable), such as AD (lask2023incontext) and DPT (lee2023supervised). In general, the above approaches rely on high-quality samples or domain knowledge to infer task beliefs during evaluation, limiting their generality with this restricted form of supervision. This inspires us to explore a much broader source of supervision from natural language.
@@
LLMs for RL. Existing studies on leveraging LLMs for RL domains can be categorized into three main threads: LLMs as rewards, LLMs as policies, and language-conditioned policy learning. The first kind uses LLMs or VLMs (vision-language models) to automate the generation and shaping of reward functions (kwon2023reward; rocamonde2024vision; xie2024text2reward; fu2024furl) or state representations (wang2024llm), providing denser information to facilitate RL training. The second kind utilizes LLMs or VLMs as the policy backbone (ahn2022can; zhai2024fine; shi2024unleashing; hu2025diversity). Vision-language-action (VLA) models provide a direct instantiation of using pre-trained VLMs for robotics, fine-tuning visually-conditioned VLMs to generate robot control actions, such as RT-2 (brohan2023rt), OpenVLA (kim2024openvla), and $\pi_{0.5}$ (intelligence2025pi_). In contrast, we leverage LLMs to obtain task representations for offline meta-RL and train conventional RL policies based on decision transformer or diffuser architectures.
@@
Language-conditioned policy. We focus on the third kind that learns a language-conditioned policy, extracting the world knowledge encoded in LLMs to help train offline meta-RL agents. Previous studies acquire language embeddings using pre-trained LLMs like BERT and GPT (hill2020human), while these embeddings are learned independently from decision tasks and may fail to capture domain-related information. Early methods develop rule-based (alomari2017natural) or task-specific (akakzia2021grounding) intermediate representations to capture task-related semantics, which can require a laborious design and lack scalability. Under an imitation learning paradigm, BC-Z (jang2022bc) trains a vision-based robotic manipulation policy conditioned on pre-trained language embeddings, and BAKU (haldar2024baku) produces a simple transformer architecture that fuses multi-modal vision-language information and temporal context for multi-task decision-making. To improve language grounding, (pang2023natural) translates natural language to task language with a referential game, and (pang2024knowledgeable) uses supervised fine-tuning to enable bidirectional translation between language-described skills and rollout data. Overall, the challenges include the inability to capture environment dynamics, knowledge misalignment due to semantic gaps, and limited model scalability. In the paper, we tackle these challenges to fully harness LLMs for building generalist decision agents.
@@ 3 Method
In this section, we present Text-to-Decision Agent (T2DA), a simple and scalable framework to train generalist policies. Figure 2 illustrates the overall pipeline, followed by detailed implementations in the subsequent subsections. Corresponding algorithm pseudocodes are given in Appendix A.
@@ 3.3 Contrastive Language-Decision Pre-training
To successfully instruct generalist training with human language, we must bridge the semantic gap between text and the “decision modality". Inspired by CLIP (radford2021learning), we predict which textual description goes with which training task, bridging their semantic gap via contrastive learning. The core idea is to connect the language supervision with corresponding decision embeddings, distilling the world model structure to the text modality and aligning text embeddings to comprehend environment dynamics.
Formally, given a batch of $N$ (trajectory $\tau$, textual task description $l$) pairs, we predict which of the $N\times N$ possible (trajectory, text) pairings across a batch actually occurred. We use the pre-trained trajectory encoder in Sec. 3.2 to extract the dynamics-aware decision embedding from the trajectory as $z=\phi(\tau)$. The text encoder is a transformer-based model $\psi$ that could be initialized from one of a wide variety of popular language models, such as CLIP (radford2021learning) or T5 (raffel2020exploring). Then, the text embedding $z_{T}$ is easily derived by tokenizing the textual instruction $l$ and feeding it to the encoder as $z_{T}=\psi(l)$.
We model the pre-training task as a text-decision multi-modal learning problem by fine-tuning the text encoder while keeping the trajectory encoder fixed. The objective is to maximize the cosine similarity of the decision and text embeddings of the $N$ real pairs in the batch while minimizing the cosine similarity of the embeddings of the $N^{2}-N$ incorrect pairings. The two directional similarities between the modalities, which are transposes of each other, are defined as
[…]
@@
$p(\tau_{k})=\frac{e^{\text{sim}(\tau_{k},l_{k})}}{\sum_{i=1}^{N}e^{\text{sim}(\tau_{k},l_{i})}},~~~p(l_{k})=\frac{e^{\text{sim}(l_{k},\tau_{k})}}{\sum_{i=1}^{N}e^{\text{sim}(l_{k},\tau_{i})}}.$ (4)
@@
where CE is the cross-entropy between two distributions. To effectively adapt the text encoder to the semantics of decision modality at a lightweight computing cost, we employ low-rank adaptation (LoRA) (hu2022lora) to fine-tune the text encoder with the above contrastive loss. Experiments in Appendix E present the analysis of its efficiency and superiority compared to full-parameter fine-tuning.
@@ 3.4 Generalist Policy Learning
A widely adopted manner to tackle generalization in RL domains is to condition the policy on some representation that can capture task-relevant information, such as context-based meta-learning (rakelly2019efficient; zintgraf2021varibad) and prompt-based methods (xu2022prompting; wang2024meta). Under this unified framework, we formulate the generalist agent as a task-conditioned policy model. There exists a true variable that represents the task identity, and we use a latent representation $h$ to approximate that variable. Then, a base policy $\pi(\cdot,h)$ can be shared across tasks by conditioning on the task representation as $\pi_{k}(a|s)\!=\!\pi(a|s;h_{k}),\forall k$. Previous studies usually approximate the task representation by acquiring expert decision data (xu2022prompting) or exploring the unseen task first (gao2023context). This restricted form of supervision limits their universality and practicability since expensive high-quality samples or explorations are needed to specify any new task.
@@
3.5 Scalable Implementations
We develop scalable implementations of T2DA using two mainstream conditional generation architectures that hold the promise to train RL models at scale: the decision diffuser (ajay2023conditional) and the decision transformer (chen2021decision), yielding the following T2DA variants.
Text-to-Decision Diffuser (T2DA-D). Inspired by the recent success of text-to-image diffusion models (saharia2022photorealistic; zhang2023adding), we implement the text-to-decision counterpart by modeling the policy as a return-conditioned diffusion model with action planning. For model simplicity and scalability, we choose to diffuse over the $H$-step state-action trajectory as
@@ 3.5 Scalable Implementations
where $c$ denotes the timestep in the forward diffusion process, and c=0c\!=\!0 corresponds to the unperturbed data.  [^1] Decision diffuser diffuses over only states and trains another inverse dynamics model to generate actions. In contrast, we only keep a unified diffusion model to i) ensure simplicity with a single end-to-end model in a multi-task setting, and ii) unlock the potential of diffusion models at scale, rather than depending on additional MLPs to generate part of the information, e.g., actions. To enable a generalist policy with natural language supervision, we include the fine-tuned text embedding $\psi(l)$ as an additional condition to the diffusion model. To this end, the generalist policy training is formulated as the standard problem of conditional generative modeling as
@@
Text-to-Decision Transformer (T2DA-T). We also implement T2DA with the causal transformer architecture, drawing upon the simplicity, scalability, and associated advances in language modeling such as BERT and GPT. We prepend a task prompt to the input of the causal transformer to realize generalization across tasks, akin to Prompt-DT (xu2022prompting). For each task, we take the prompt-augmented trajectory $\tau^{+}$ as the input, which contains both the fine-tuned text embeddings $\psi(l)$ and the most recent $H$-step history sampled from offline datasets as
@@ 4 Experiments
Environments. We evaluate T2DA on three benchmarks that are widely adopted to assess generalization capacities of RL algorithms: i) the 2D navigation Point-Robot; ii) the multi-task MuJoCo locomotion control, containing Cheetah-Vel and Ant-Dir; and iii) the Meta-World platform for robotic manipulation, where a robotic arm is designed to perform a wide range of manipulation tasks, such as close faucet, lock door, open door, and press button. For each domain, we randomly sample a distribution of tasks captioned with text descriptions, and split them into training and test sets. We employ SAC (haarnoja2018soft) to independently train a single-task policy on each training task for offline dataset collection. We develop three types of offline datasets: Mixed, Medium, and Expert, following the common practice in offline RL (fu2020d4rl). Appendix B presents more details of environments and datasets.
Baselines. We compare T2DA to five competitive baselines that cover three representative paradigms in tackling RL generalization: the context-based offline meta-RL approaches, including 1) CSRO (gao2023context) and 2) UNICORN (li2024towards); the in-context RL method of 3) AD (lask2023incontext); and the language-conditioned policy learning methods including 4) BC-Z (jang2022bc) and 5) BAKU (haldar2024baku).
For evaluation, all experiments are conducted using five different random seeds. The mean of received return is plotted as the bold line, with 95% bootstrapped confidence intervals of the mean indicated by the shaded region. Appendix D gives detailed implementations of T2DA. Appendix E gives analysis on fine-tuning of the text encoder.